A car traveled 100 kilomters in 2 hours. What is the average speed the car traveled? show your work.
Answer by checkley77(12844)

DISTANCE=RATE*TIME
D=RT
R=D/T
R=100/2=50 KPH. IS THE SPEED (RATE).